OpenStudy (anonymous):

Two spheres have a volume ratio of 27:8. What is the ratio of their radii?

OpenStudy (anonymous):

I don't get how to do this.

OpenStudy (anonymous):

@JoannaBlackwelder

OpenStudy (joannablackwelder):

I think that the radii would be the ratio of the cube roots of the volumes.

OpenStudy (anonymous):

Let say the volume of big sphere is V1 and the volume of small one is V2, and we have \(\dfrac{V1}{V2}=\dfrac{27}{8}\) ok so far? not finish yet, but I need know you got it or not.

OpenStudy (anonymous):

@claritamontano I am waiting for your reply to step up.

OpenStudy (anonymous):

Yeah I get it,

OpenStudy (anonymous):

We have formula to find the volume of the sphere. It says V= \(\dfrac{4}{3}\pi r^3\) where r is radius of the sphere, so that, we have \(V1= \dfrac{4}{3}\pi r_1^3\) and \(V2= \dfrac{4}{3}\pi r_2^3\) so far so good?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

so, just plug in V1/V2 = 27/8 you have \(\dfrac {r_1^3}{r_2^2}=\dfrac{27}{8}\) or \(\dfrac{r_1}{r_2}= \dfrac{3}{2}\)

OpenStudy (anonymous):

ah I get it so my answer would now be 3:2

OpenStudy (anonymous):

got it? just algebra \(\dfrac {r_1^3}{r_2^2}=(\dfrac{r_1}{r_2})^3=\dfrac{27}{8}=(\dfrac{3}{2})^3\) which gives us the ratio of the radii is 3/2

OpenStudy (anonymous):

yeah! thanks wait not 3:2 or is that the same thing?

OpenStudy (anonymous):

the same 3:2

OpenStudy (anonymous):

okay thanks I get it!